solve system using addition method
x+2y=20
5x-2y+16
solution
(?,?)

Answer by checkley77(12844)

x+2y=20
5x-2y[=]16 add. I'll assume this is an = sign.
------------
6x=36
x=36/6
x=6 ans.
6+2y=20
2y=20-6
2y=14
y=14/2
y=7 ans.